# LLM Maybe LongLM: SelfExtend LLM Context Window Without Tuning

### v2

@@ 1 Introduction
Recently, a variety of context window extension methods have been developed to tackle the challenge of extending the context window size of pretrained LLMs. A straightforward approach is to fine-tune these models on enough extensive texts. Besides this, some methods seek to extend context window length in more efficient fine-tuning ways. Among these contemporary methods, some notable techniques include ‘PI’ (Chen et al., 2023b), ‘CLEX’ (Chen et al., 2023a) ‘Yarn’ (Peng et al., 2023), ‘LongLora’ (Chen et al., 2023c), and ‘ABF’ (Xiong et al., 2023). These methods aim to extend the content window based on the implicit assumption that pretrained LLMs lack the ability to handle long content. However, these methods typically require finetuning to achieve extension, which can be resource and time intensive given the quadratic complexity of Transformers. Additionally, high-quality long text data is scarce, hindering such fine-tuning approaches. Most real-world data is short, and much long text lacks meaningful long-range dependencies. With limited appropriate data, finetuning risks degrading existing strong performance on shorter sequences from pretraining or overfitting models to the tuning set. LLMs’ generalizability to broad tasks may reduce.
Instead of extending the content window, in this paper, we believe LLMs should have inherent capabilities to handle long contexts. Our belief stems from the fact that when we, as human beings, are children, we are taught how to read and write using relatively short texts, such as articles spanning several pages. We rarely use extremely long texts like entire books or complete documents as learning materials. Yet, we are still able to understand long texts effectively. With this strong motivation, the poor performance of LLMs while facing long text out of the pretraining context window is not due to the lack of long context understanding capabilities.
@@
In our analysis, the key challenge preventing LLMs from effectively handling longer contexts is the Out-of-Distribution (O.O.D) issues related to positional encoding, which we call the positional O.O.D [^1] Here, the position refers to relative position rather than absolute position. The relative position is $m-n$ in RoPE, where $m$ and $n$ are the absolute positions of two tokens. The positional O.O.D refers to cases where the value of $m-n$ during inference is unseen, i.e., larger than the values observed during pretraining. In this paper, we map unseen large relative positions to those observed during pretraining. More details about $m-n$ are provided in Section 2. issue. This problem arises when LLMs encounter text sequences during inference exceeding the length of their pretraining context window, where LLMs are exposed to new relative distances that were not present during their pretraining phase. It is widely recognized that Neural Networks (NNs) are susceptible to unpredictable behaviors when dealing with O.O.D inputs (Liu et al., 2021; Shen et al., 2021; Bai et al., 2021; Zhang et al., 2023). To address this, an intuitive and practical solution would be to remap the unseen relative positions to those encountered during the pretraining, thus extending the LLMs’ ability to handle longer contexts naturally.
@@
Our SelfExtend is a plug-and-play method that takes effect at the inference stage, allowing existing large language models to easily adopt it. We evaluate SelfExtend with some popular LLMs (Llama-2 (Touvron et al., 2023), Mistral (Jiang et al., 2023), SOLAR (Kim et al., 2023), and Phi-2 (Javaheripi et al., 2023)) on three types of tasks: language modeling, synthetic long context tasks, and real-world long context tasks. The proposed SelfExtend substantially improves the long context understanding ability and even outperforms many finetuning-based methods on some tasks. These results underscore SelfExtend as an effective solution for context window extension. The superior performance of SelfExtend also demonstrated the potential of large language models to effectively handle long contexts. Our main contributions are summarized as follows:
@@ 2 Preliminary
Position Encoding. Transformers (Vaswani et al., 2017) incorporate position information via different positional embedding designs. The positional embedding design can be categorized into two classes: absolute position embeddings and relative positional encodings. The absolute position embedding provides the absolute positions, which embeds each absolute position $i$ into position vector $\mathbf{p}_{i}$ and adds word embeddings to their corresponding $\mathbf{p}_{i}$ before feeding them to the model. Examples of such include sinusoidal position embeddings (Vaswani et al., 2017) and learned position embeddings in GPT3 (Brown et al., 2020) and OPT (Zhang et al., 2022), or adding the dot product between two tokens’ position embeddings on the attention logit (Ke et al., 2020).
On the other hand, relative positional encodings have been proposed to use relative distance information between tokens and have become the mainstream of position embedding. This information is usually applied in attention layers. Examples of such include a learnable attention logit bias as in T5 (Xue et al., 2020), Transformer-XL (Dai et al., 2019); a fixed linear attention decay called Alibi (Press et al., 2021); rotating query and key sequences based on distance such as RoPE (Su et al., 2022), and XPos (Sun et al., 2023). The proposed method is based on the Rotary Position Embedding (RoPE) introduced in (Su et al., 2022).
@@ 3.1 Preliminary Analysis
① Why do LLMs fail on sequences during inference that are longer than their pre-training context window? For a pretrained LLM with relative position encodings, such as RoPE, the behavior of the LLMs becomes unpredictable during inference if the length of a sequence is longer than its pretraining context window length. This has been explored by (Han et al., 2023; Chen et al., 2023b) that with unseen relative positions, the attention distributions are very different compared to those within the pretraining context window length. We argue that such failure stems from the Out-of-Distribution (O.O.D.) relative distance in the same sense that neural networks are not robust to O.O.D. inputs (Shen et al., 2021).
@@
④ How to restore degraded language modeling ability caused by grouped attention? — Re-introducing normal attention in the neighboring area. In the process of generating next tokens, the immediate neighbors of a target token play a crucial role, which is well-supported by existing methods of sparse attention mechanisms (Zaheer et al., 2020; Shi et al., 2021) and methods for extending the contextual window (Han et al., 2023; Xiong et al., 2023; Chen et al., 2023c). These studies consistently highlight the importance of maintaining the standard attention mechanism for tokens in close proximity to the target token. This proximity-based focus is essential for the accurate generation of the next token, ensuring the coherence and fluency of the generated text, as evidenced by acceptable perplexity (PPL) levels. Employing grouped attention might not significantly affect the overall quality of generated sentences; however, it necessitates the accurate positioning of attention to maintain generation quality. Therefore, it is imperative to preserve the standard attention mechanism within the vicinity of the target token, as utilized during the pretraining phase, to ensure the precision and effectiveness of language models in capturing the nuances of local context.
@@ 3.2 SelfExtend LLM Context Window Without Tuning
For example, in Figure 3, the context window is extended from its pretraining length of $7$ to $(7-4)*2+4=10$. The pseudo code for SelfExtend are presented in Algorithm 1.
@@ 4 Experiments
We evaluate SelfExtend with Llama-2 (Touvron et al., 2023) and its families, Phi-2 (Javaheripi et al., 2023), Mistral (Jiang et al., 2023) and SOLAR (Kim et al., 2023) on language modeling task, synthetic long context tasks, real-world long context tasks and standard short-context tasks.
@@ 4.1 Performance on Language Modeling Tasks
Language modeling task is the most fundamental and the least requirement for LLMs, which is usually measured by perplexity (PPL) on the test text data. A low PPL does not guarantee good performance on real tasks (Pal et al., 2023), however, a higher PPL suggests severe performance degradation of LLMs. We evaluate SelfExtend’s language modeling performance on dataset PG19 (Rae et al., 2019), which contains lengthy books. PPL is used as the metric. More experimental details are presented in Section D.1
The results show that SelfExtend can successfully maintain a low PPL out of the pretraining context window for both Llama-2-7b-chat and Mistral. Without SelfExtend, the PPL explodes when the length of test sequence is larger than the context window. Mistral with SWA can also maintain a low PPL out of its context window. But later in the next section, we will demonstrate that a low PPL score does not necessarily indicate proficiency in handling long contexts. More discussion about PPL can be found in Appendix B.
@@ 4.3 Performance on Real-World Long Context Tasks
Evaluation solely on language modeling (measured by perplexity) and synthetic tasks like passkey retrieval cannot fully assess the long-context capabilities of LLMs. The task of Passkey retrieval is overly straightforward, and an LLM may still struggle with long context despite low perplexity. To comprehensively evaluate long-context performance, we further use two recent real-world long context benchmarks: LongBench (Bai et al., 2023) and L-Eval (An et al., 2023). The results are presented in Table 2 and Table 3.
On the LongBench in Table 2, for all four different base LLMs and most datasets, with SelfExtend, the LLM can obtain significant performance improvments.
@@
Mistral-7B: We extend Mistral-7B’s context to 16k, significantly improving its long context ability over the base model, with or without SWA applied. The fine-tuned variant MistralLite ((amazon, 2023)) achieves the best performance on most datasets. However, many of these datasets were included in MistralLite’s fine-tuning data, such as NarrativeQA [^2] More details about MistralLite’s fine-tuning data can be found at https://huggingface.co/amazon/MistralLite. At least, GovReport, QMSum, NarrativeQA, Qasper, QuALITY, and HotpotQA are included. Meanwhile, Multi-passage QA and summarization tasks are also in fine-tuning data. This also violates zero-shot evaluation conditions..
@@ 5 Conclusion and Discussion
In this paper, we argue that LLMs themselves have the inherent ability to handle long sequences and propose SelfExtend to elicit the inherent long context abilities for LLMs by simply mapping unseen relative positions into those seen during pretraining via the Floor operation. Without any tuning or further training, SelfExtend can effectively improve LLMs’ long context performance, as extensive experiments show.
Limitations: SelfExtend increases computation cost with naive implementations since it performs extra attention across all query-key pairs. However, with optimizations like blocked kernels (e.g., Flash Attention (Dao et al., 2022)), this becomes linear rather than quadratic, and the marginal cost is small enough to be ignored for long input sequences. Also, the performance degrades with large group size, preventing indefinitely long contexts. Additionally, evaluation methodologies for assessing long context abilities remain open research questions.
Future Work: We are interested in testing SelfExtend on models using other positional encoding. Larger models, longer contexts, and more challenging tasks will be tested if we can access more computational resources in the future. In the meantime, more sophisticated mapping methods will be considered as the replacement of the simple floor operation to achieve better long context understanding abilities and extended context window length.
@@ Appendix B Perplexity as a Metric for Long Context Capabilities
PPL is not an effective metric for measuring the ability of LLMs to handle long contexts. In Figure 7, we introduce a seemingly plausible context window extension method named ’Infinite’. When evaluated on PG19 using the same protocol, Llama-2-7b-chat with ’Infinite’ achieves PPL scores that are comparable to, or even lower than, those achieved by SelfExtend, as demonstrated in Table 6. However, ’Infinite’ essentially mimics the process of dividing a long sequence into short sub-sequences before processing them with LLMs, indicating that it does not genuinely address long context handling.
@@
Table 6: Perplexity on the PG19 dataset: For ’Infinite’, we set three different local window sizes: 1024, 2048, and 4096. We have also included the results from Table 1 for comparison.
@@
Appendix C SelfExtend on Group Size and Neighbor Window
@@ D.3 Experimental Setting on Varying-Length Passkey Retrieval Task
In this experiment, we use the following models: Llama2-7b-chat with SelfExtend, LongLora-7b-16k [^6] We use its fully fine-tuned variant. For more details about the model: https://huggingface.co/Yukang/Llama-2-7b-longlora-16k. We didn’t use the latest version named ’LongAlpaca’ as we cannot get reasonable performance for this specific task with LongAlpaca, which may be due to our improper configuration.,vicuna-1.5-7b-16k, Together AI’s Llama-2-7b-32k [^7] Both vicuna-1.5-7b-16k and Together AI’s Llama-2-7b-32k were fine-tuned using position interpolation, and Yarn-Llama-2-7b-64k.
Appendix E Details of LLMs
@@
Model Name
URL
Llama-2-7b-chat-hf (Touvron et al., 2023)
https://huggingface.co/meta-llama/Llama-2-7b-chat-hf
Mistral-7B-Instruct-v0.1 (Jiang et al., 2023)
https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.1
Phi-2 (Javaheripi et al., 2023)
https://huggingface.co/microsoft/phi-2
SOLAR-10.7B-Instruct-v1.0 (Kim et al., 2023)
https://huggingface.co/upstage/SOLAR-10.7B-Instruct-v1.0
LongChat-7b-v1.5-32k
https://huggingface.co/lmsys/longchat-7b-v1.5-32k
togethercomputer/LLaMA-2-7B-32K
https://huggingface.co/togethercomputer/LLaMA-2-7B-32K
CLEX-7B-16K (Chen et al., 2023a)
https://huggingface.co/DAMO-NLP-SG/CLEX-7B-16K
CodeLlama-7b-hf (Rozière et al., 2023)
https://huggingface.co/codellama/CodeLlama-7b-hf
vicuna-7b-v1.5-16k
https://huggingface.co/lmsys/vicuna-7b-v1.5-16k
MistralLite (amazon, 2023)
https://huggingface.co/amazon/MistralLite

### v3

@@ 1 Introduction
Recently, a variety of context window extension methods have been developed to extend the context window of pretrained LLMs. A straightforward approach is to fine-tune these models on enough extensive texts. Besides this, some methods seek to extend context window length in more efficient fine-tuning ways. Among these contemporary methods, some notable techniques include ‘PI’ (Chen et al., 2023b), ‘CLEX’ (Chen et al., 2023a) ‘Yarn’ (Peng et al., 2023), ‘PoSE’ (Zhu et al., 2023), ‘LongLora’ (Chen et al., 2023c), and ‘ABF’ (Xiong et al., 2023). These methods aim to extend the content window based on the implicit assumption that pretrained LLMs lack the ability to handle long content. However, these methods typically require finetuning to achieve extension, which can be resource and time-intensive given the quadratic complexity of Transformers. Additionally, high-quality long text data is scarce, hindering such fine-tuning approaches. Most real-world data is short, and much long text lacks meaningful long-range dependencies. With limited appropriate data, finetuning risks degrading existing strong performance on shorter sequences from pretraining or overfitting models to the tuning set. LLMs’ generalizability to broad tasks may be reduced.
@@
Instead of extending the content window, in this paper, we believe LLMs should have inherent capabilities to handle long contexts. Our belief stems from the fact that when we, as human beings, are children, we are taught how to read and write using relatively short texts, such as articles spanning several pages. We rarely use extremely long texts like entire books or complete documents as learning materials. Yet, we are still able to understand long texts effectively. With this strong motivation, the poor performance of LLMs while facing long text is not due to the lack of long context understanding capabilities. In our analysis, the key challenge preventing LLMs from effectively handling longer contexts is the Out-of-Distribution (O.O.D) issues related to positional encoding, which we call the positional O.O.D [^1] Here, the position refers to relative position rather than absolute position. The relative position is $m-n$ in RoPE, where $m$ and $n$ are the absolute positions of two tokens. The positional O.O.D refers to cases where the value of $m-n$ during inference is unseen, i.e., larger than the values observed during pretraining. In this paper, we map unseen large relative positions to those observed during pretraining. More details about $m-n$ are provided in Section 2. issue. This problem arises when LLMs encounter text sequences during inference exceeding the length of their pretraining context window, where LLMs are exposed to new relative distances that were not present during their pretraining phase. It is widely recognized that Neural Networks (NNs) are susceptible to unpredictable behaviors when dealing with O.O.D inputs (Liu et al., 2021; Shen et al., 2021; Bai et al., 2021; Zhang et al., 2023). To address this, an intuitive and practical solution would be to remap the unseen relative positions to those encountered during the pretraining, thus extending the LLMs’ ability to handle longer contexts naturally.
@@
Our SelfExtend is a plug-and-play method that takes effect at the inference stage, allowing existing large language models to easily adopt it. We evaluate SelfExtend with some popular LLMs (Llama-2 (Touvron et al., 2023), Mistral (Jiang et al., 2023a), SOLAR (Kim et al., 2023), and Phi-2 (Javaheripi et al., 2023)) on three types of tasks: language modeling, synthetic long context tasks, and real-world long context tasks. The proposed SelfExtend substantially improves the long context understanding ability and even outperforms many finetuning-based methods on some tasks. These results underscore SelfExtend as an effective solution for context window extension. The superior performance of SelfExtend also demonstrated the potential of large language models to effectively handle long contexts. Our main contributions are summarized as follows:
@@ 2 Preliminary
In this section, we present the preliminaries of our work.
Position Encoding. Transformers (Vaswani et al., 2017) incorporate position information via different positional embedding designs. The common positional embedding design can generally be categorized into two classes: absolute position embeddings and relative positional encodings. The absolute position embedding provides the absolute positions, which embeds each absolute position $i$ into position vector $\mathbf{p}_{i}$ and adds word embeddings to their corresponding $\mathbf{p}_{i}$ before feeding them to the model. Examples of such include sinusoidal position embeddings (Vaswani et al., 2017) and learned position embeddings in GPT3 (Brown et al., 2020) and OPT (Zhang et al., 2022), or adding the dot product between two tokens’ position embeddings on the attention logit (Ke et al., 2020). On the other hand, relative positional encodings have been proposed to use relative distance information between tokens and have become the mainstream of position embedding. This information is usually applied in attention layers. Examples of such include a learnable attention logit bias as in T5 (Xue et al., 2020), Transformer-XL (Dai et al., 2019); a fixed linear attention decay called Alibi (Press et al., 2021); rotating query and key sequences based on distance such as RoPE (Su et al., 2022), and XPos (Sun et al., 2023). The proposed method in this work is based on the Rotary Position Embedding (RoPE) introduced in (Su et al., 2022).
@@ 3.1 Preliminary Analysis
① Why do LLMs fail on sequences during inference that are longer than their pre-training context window? For a pretrained LLM with relative position encodings, such as RoPE, the behavior of the LLMs becomes unpredictable during inference if the length of a sequence is longer than its pretraining context window length. This has been explored by (Han et al., 2023; Chen et al., 2023b) that with unseen relative positions, the attention distributions are different compared to those within the pretraining context window. We argue that such failure stems from the Out-of-Distribution (O.O.D.) relative distance in the sense that neural networks are not robust to O.O.D. inputs (Shen et al., 2021).
@@
④ How to restore degraded language modeling ability caused by grouped attention? — Re-introducing normal attention in the neighboring area. In the process of generating next tokens, the immediate neighbors of a target token play a crucial role, which is well-supported by existing methods of sparse attention mechanisms (Zaheer et al., 2020; Shi et al., 2021) and methods for extending the contextual window (Han et al., 2023; Xiong et al., 2023; Liu et al., 2024). These studies consistently highlight the importance of maintaining the standard attention mechanism for tokens in close proximity to the target token. This proximity-based focus is essential for the accurate generation of the next token, ensuring the coherence and fluency of the generated text, as evidenced by acceptable perplexity (PPL) levels. Employing grouped attention might not significantly affect the overall quality of generated sentences; however, it necessitates the accurate positioning of attention to maintain generation quality. Therefore, it is imperative to preserve the standard attention mechanism within the vicinity of the target token, as utilized during the pretraining phase, to ensure the precision and effectiveness of language models in capturing the nuances of local context.
@@ 3.2 SelfExtend LLM Context Window Without Tuning
For example, in Figure 3, the context window is extended from its pretraining length of $7$ to $(7-4)*2+4=10$. The pseudo code for SelfExtend is presented in Algorithm 1.
@@ 4 Experiments
We evaluate SelfExtend with Llama-2 (Touvron et al., 2023) and its families, Phi-2 (Javaheripi et al., 2023), Mistral (Jiang et al., 2023a) and SOLAR (Kim et al., 2023) on language modeling task, synthetic long context tasks, real-world long context tasks and standard short-context tasks.
@@
Gradient-Llama-3-8B-Inst-262k(32k) + 21.71 44.24 44.54 46.82 36.42 21.49 30.03 22.67 27.79 74.5 90.23 42.53 NA 67.00 57.00 51.22
Gradient-Llama-3-8B-Inst-1M(32k) + 12.04 12.10 20.15 8.22 9.68 3.89 27.90 14.58 22.13 61.00 82.82 1.40 2.37 2.83 57.87 56.42
SE-Llama-3-8B-ins-32k    10/96 + 12.04 12.10 20.15 8.22 9.68 3.89 27.90 14.58 22.13 61.00 82.82 1.40 2.37 2.83 57.87 56.42
@@ 4.1 Performance on Language Modeling Tasks
Language modeling task is the most fundamental and the least requirement for LLMs, which is usually measured by perplexity (PPL) on the test text data. A low PPL does not guarantee good performance on real tasks (Pal et al., 2023), however, a higher PPL suggests severe performance degradation of LLMs. We evaluate SelfExtend’s language modeling performance on dataset PG19 (Rae et al., 2019), which contains lengthy books. PPL is used as the metric. More experimental details are presented in Section D.1 The results show that SelfExtend can successfully maintain a low PPL out of the pretraining context window for both Llama-2-7b-chat and Mistral. Without SelfExtend, the PPL explodes when the length of test sequence is larger than the context window. Mistral with SWA can also maintain a low PPL out of its context window. But later in the next section, we will demonstrate that a low PPL score does not necessarily indicate proficiency in handling long contexts. More discussion about PPL can be found in Appendix B.
@@ 4.3 Performance on Real-World Long Context Tasks
Evaluation solely on language modeling (measured by perplexity) and synthetic tasks like passkey retrieval cannot fully assess the long-context capabilities of LLMs. The task of Passkey retrieval is overly straightforward, and an LLM may still struggle with long context despite low perplexity. To comprehensively evaluate long-context performance, we further use two recent real-world long context benchmarks: LongBench (Bai et al., 2023) and L-Eval (An et al., 2023). The results are presented in Table 2 and Table 3. On the LongBench in Table 2, for all four different base LLMs and most datasets, with SelfExtend, the LLM can obtain significant performance improvements.
@@
Mistral-7B: We extend Mistral-7B’s context to 16k, significantly improving its long context ability over the base model. The fine-tuned variant MistralLite ((amazon, 2023)) achieves the best performance on most datasets. However, many of these datasets were included in MistralLite’s fine-tuning data, such as NarrativeQA [^2] More details about MistralLite’s fine-tuning data can be found at https://huggingface.co/amazon/MistralLite. At least, GovReport, QMSum, NarrativeQA, Qasper, QuALITY, and HotpotQA are included. Meanwhile, Multi-passage QA and summarization tasks are also in fine-tuning data. This also violates zero-shot evaluation conditions..
@@ 4.5 Ablations on Group Size and Neighbor Window
More results can be found in Appendix G. When the group size is so large that all distant tokens are in one group, SelfExtend degenerates into ReRoPE (e.g. group size of 32768 in Figure 10).
@@ 5 Conclusion and Discussion
In this paper, we argue that LLMs themselves have the inherent ability to handle long sequences and propose SelfExtend to elicit the inherent long context abilities for LLMs by mapping unseen relative positions into those seen during pretraining. Without any tuning or further training, SelfExtend can effectively improve LLMs’ long context performance.
Limitations: SelfExtend increases computation cost with naive implementations, since it performs extra attention across all query-key pairs. However, with optimizations like blocked kernels (e.g. Flash Attention (Dao et al., 2022)), this becomes linear rather than quadratic, and the marginal cost is small enough to be ignored for long input sequences. Also, the performance degrades with large group size, preventing indefinitely long contexts. Besides, SelfExtend still processes the entire sequence to ensure information integrity, while some methods such as prompt compression (Chuang et al., 2024; Jiang et al., 2023b) can shorten the input to reduce computation. Additionally, evaluation methodologies for assessing long context abilities remain open research questions. Standard practices have yet to emerge, complicating experimental results.
Future Work: We are interested in more sophisticated mapping methods to replace the simple floor operation, aiming to enhance long context understanding and extend the context window length. Additionally, we plan to further investigate the complex behaviors of LLMs using SelfExtend.
Impact Statement
The impact of this contribution is multifaceted. Firstly, it enhances the usability and applicability of LLMs across various domains that require the processing of lengthy input sequences, such as document analysis, long-form question answering, and retrieval augmented generation. Secondly, the ability to extend the context window without fine-tuning simplifies the deployment of advanced language models, making them more accessible to a broader range of users and applications. The availability of the code makes this solution readily accessible to researchers and practitioners, promising widespread adoption and further innovation in the field of natural language processing.
Acknowledgements
The authors thank the anonymous reviewers for their helpful comments. This work is in part supported by NSF grants NSF IIS-2310260 and IIS-2224843. The views and conclusions contained in this paper are those of the authors and should not be interpreted as representing any funding agencies.
@@ Appendix B Perplexity as a Metric for Long Context Capabilities
PPL is not an effective metric for measuring the ability of LLMs to handle long contexts. In Figure 7, we introduce a seeming plausible context window extension method named ’Infinite’. When evaluated on PG19 using the same protocol, Llama-2-7b-chat with ‘Infinite’ achieves PPL scores that are comparable to, or even lower than, those achieved by SelfExtend, as demonstrated in Table 6. However, ‘Infinite’ essentially mimics the process of dividing a long sequence into short sub-sequences before processing them with LLMs, indicating that it does not genuinely address long context handling.
@@
Table 6: Perplexity on the PG19 dataset: For ‘Infinite’, we set three different local window sizes: 1024, 2048, and 4096. We have also included the results from Table 1 for comparison.
@@
Appendix C SelfExtend with Varying Group Size and Neighbor Window
@@ D.3 Experimental Setting on Varying-Length Passkey Retrieval Task
In this experiment, we use the following models: Llama2-7b-chat with SelfExtend, LongLora-7b-16k [^6] We use its fully fine-tuned variant, as we cannot use the LongAlpaca version to get reasonable performance for this specific task. For more details about the model: https://huggingface.co/Yukang/Llama-2-7b-longlora-16k,vicuna-1.5-7b-16k, Together AI’s Llama-2-7b-32k [^7] Both vicuna-1.5-7b-16k and Together AI’s Llama-2-7b-32k were fine-tuned using position interpolation, and Yarn-Llama-2-7b-64k.
Appendix E Detail of LLMs
@@
Model Name URL
Llama-2-7b-chat-hf (Touvron et al., 2023) https://huggingface.co/meta-llama/Llama-2-7b-chat-hf
Mistral-7B-Instruct-v0.1 (Jiang et al., 2023a) https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.1
Phi-2 (Javaheripi et al., 2023) https://huggingface.co/microsoft/phi-2
SOLAR-10.7B-Instruct-v1.0 (Kim et al., 2023) https://huggingface.co/upstage/SOLAR-10.7B-Instruct-v1.0
LongChat-7b-v1.5-32k https://huggingface.co/lmsys/longchat-7b-v1.5-32k
togethercomputer/LLaMA-2-7B-32K https://huggingface.co/togethercomputer/LLaMA-2-7B-32K
CLEX-7B-16K (Chen et al., 2023a) https://huggingface.co/DAMO-NLP-SG/CLEX-7B-16K
CodeLlama-7b-hf (Rozière et al., 2023) https://huggingface.co/codellama/CodeLlama-7b-hf
vicuna-7b-v1.5-16k https://huggingface.co/lmsys/vicuna-7b-v1.5-16k
MistralLite (amazon, 2023) https://huggingface.co/amazon/MistralLite
Appendix F SelfExtend with Other Positional Encodings
In this section, we test SelfExtend on LLMs using non-RoPE positional encodings. We implemented SelfExtend for MPT-7b-chat (Team, 2023), which uses Alibi (Press et al., 2021) as its positional embedding method. We conducted experiments on the PG19 dataset. The results are shown in Table 8. The results show that SelfExtend is able to work with non-RoPE positional encodings, which are pretty similar to models with RoPE positional encoding,
Table 8: Perplexity of MPT-7b-chat on PG19 with different sequence length. The vanilla MPT-7b-chat has a context window of 2k tokens. For SelfExtend, We set the neighbor window as 512 and set the group size as 6.
MPT-7b-chat 1024 2048 3172 4096 5120 6144 8192
Vanilla (2k) 8.8 9.6 12.0 26.3 52.0 115.5 196.0
SelfExtend 8.9 9.9 10.6 10.8 11.1 11.3 11.6
Appendix G Hyperparameyer Selection for SelfExtend
We conduct experiments on “Needle in a Haystack” (gkamradt, 2023), to investigate the impacts of group size and neighbor window size. The results are shown in Figure 10.
The experimental results indicate that SelfExtend is not overly sensitive to hyperparameter selection. Predefined, heuristic values for group size and neighbor window size are often sufficient to achieve satisfactory performance, as long as group size and neighbor window are not too large or too small. We conclude those results as an empirical rule. Denoting the pretraining context window as $L$, the target extension length as $N$, the neighbor window as $W$, and the group size as $G$, the empirical rule for selecting hyperparameters is to ensure that the following inequality holds:
$\frac{1}{2}\times L>W+\frac{N-W}{G}$ (5)
We believe this empirical rule is due the fact that: large relative positions are not well trained. Empirically, only a portion($\sim\frac{1}{2}$) of positions are well-trained and SelfExtend should only leverage these well-trained relative positions for the extension. This finding explains: excessively small group sizes can degrade performance, as they provide precise position information but require SelfExtend to utilize less well-trained relative positions for extension; excessively large neighbor window sizes can also degrade performance, as they provide more neighbor information but necessitate the use of less well-trained relative positions for extension.
However, the current observation may not be applicable to all models. For example, we’ve found that Llama3 series should use a much smaller neighbor window ($\sim 100$). We may dive deeper to investigate the interaction among those hyperparameters and models. Besides following the empirical rule. One could use a simple and easy-to-run representative task to find proper hyperparameters.
Figure 10: Impacts of group size and neighbor window size